OpenStudy (ammarah):

Condense the expression. Log 40 +2log4+logx

OpenStudy (anonymous):

40 tree logs + 2 tree logs + 4 tree logs + x = answr easy

OpenStudy (ammarah):

@ehuman

OpenStudy (ammarah):

is it log 46+log x?

OpenStudy (amoodarya):

log a +log b =log(a*b)

OpenStudy (amoodarya):

\[Log 40 +2\log4+logx =Log 40 +\log 4^{2}+logx \]

OpenStudy (amoodarya):

\[Log 40 +\log 4^{2}+logx =\log (40) +\log 16 +\log x\]

OpenStudy (amoodarya):

\[Log 40 \times16+logx =\log40 \times 16 \times x\]